a metal x form a salt xso4 form a clean solutions in water which react with Naoh solution so form blue precipitate y .metal x is used making electric wire and alloy like brass.what do think metal x could be?
1
Expert's answer
2013-07-01T11:38:22-0400
The metal is Copper. The salt is CuSO4. It formed whenCopper reacts with concentrated Sulfuric acid, according to the following equation:

Cu + 2H2SO4 = CuSO4 + SO2 + 2H2O

The solution of copper sulfate in water is clean and has blue color. When we add NaOH solution to CuSO4 solution, the blue precipitate forms, according to the following reaction:

CuSO4 + 2NaOH = Cu(OH)2 + Na2SO4

A Copper is used to make electric wire and brass. Hence,
x is Copper (Cu)
y is Cu(OH)2